Hyphenation Exception Log

Below is a list of words that TEX fails to hyphenate properly. It last appeared in Volume 5, No. 1, on page 15. Everything listed there is repeated here.

As expected, there are more instances of missed than of incorrect hyphens. The “TEX” column gives results from \showhyphens{...}; hyphens. \showhyphens{...} produced the results in the “TEX” column. entries in the second column are suitable for inclusion in a \hyphenation{...} list.

In most instances, inflected forms are not shown for nouns and verbs; however, all forms should be specified in a \hyphenation{...} list if they might occur in your document.

TEX

anal-yse  an-a-lyse
anomaly  anom-aly
anti-nomy(ies)  an-tin-o-my(ies)
ap-pendix  ap-pen-dix

ban-dleader  band-leader

be-haviour  be-hav-iour
bornolog-i-cal  bor-no-log-i-cal

Brow-n-ian  Brown-ian
buz-zword  buzz-word

cartwheel  cart-wheel
cholesteric  cho-les-teric
database  data-base

dat-a-p-ath  data-path
de-mos  demos

dis-tribute  dis-trib-ute

Di-jk-stra  Dijk-stra
elec-trome-chan-i-cal electro-mechan-i-cal

elec-tromechanoa-cous-tic

electro-mechano-acoustic

equiv-ari-ant  equi-vari-ant
Eu-le-rian  Euler-ian

fermions  fermi-ons

flowchart  flow-chart
Gaus-sian  Gauss-ian
ge-o-met-ric  geo-met-ric
Greif-swald  Greifs-wald

Grothendieck  Grothen-dieck
Grundlehren  Grund-leh-ren

Hamil-to-nian  Hamil-ton-ian
Her-mi-tian  Her-mit-ian
in-fras-truc-ture  in-fra-struc-ture

jeremi-ads  je-re-mi-ads

Kadomt-sev  Kadomt-sev
Leg-en-dre  Le-gendre
Lip-s-chitz(ian)  Lip-schitz(-ian)
macroe-co-nomic  macro-eco-nomic

manuscript  man-u-script

Marko-vian  Markov-ian
met-a-lan-guage  meta-lan-guage
mi-crofiche  mi-cro-fiche

mod-elling  mod-el-ling

mo-noen-er-getic  mono-en-er-getic

monopole  mono-pole

monos-trofic  mono-strofic
mul-ti-pli-ca-ble  mul-ti-plic-able
ne-ofields  neo-fields
Noethe-rian  Noether-ian
none-mer-gency  non-emer-gency

nonequiv-ari-ance  non-equi-vari-ance
noneu-clidean  non-euclid-ean

non-s-mooth  non-smooth

parametrized  pa-ram-e-trized
paramil-i-tary  para-mil-i-tary
Poincare  Poin-care
polyene  poly-ene
poly-go-niza-tion  polyg-on-i-za-tion
postam-ble  post-am-ble
pream-ble  pre-am-ble

pseu-dod-if-fer-en-tial

pseu-do-dif-fer-en-tial

pseud-ofi-nite  pseu-do-fi-nite

pseud-ofinitely  pseu-do-fi-nite-ly

pseud-o-forces  pseu-do-forces

pseu-doword  pseu-do-word
quadrat-ics  qua-drat-ics
quasiequiv-a-lence  qua-si-equiv-a-lence
quasi-hy-ponor-mal  qua-si-hy-po-nor-mal

qua-sis-mooth  qua-si-smooth
qua-sis-ta-tion-ary  qua-si-sta-tion-ary
Rie-man-nian  Rie-mann-ian

schedul-ing  sched-ul-ing
Schrodinger  Schro-ding-er
Schwarzschild  Schwarz-schild
semidef-i-nite  semi-def-in-ite
semi-ho-mo-th-etic  semi-ho-mo-thet-ies

ser-vomech-a-nism  ser-vo-mech-anism

setup  set-up
solenoid  so-le-noid
spheroid  spher-oid
stochas-tic  sto-chas-tic
sub-scriber  sub-scrib-er
summable  sum-ma-ble

ther-moe-las-tic  ther-mo-elas-tic
times-tamp  time-stamp
ve-r-all-ge-mein-erte ver-all-ge-mein-erte

wahrschein-lichkeit-s-the-o-rie

wahr-schein-lich-keits-the-o-rie

waveg-uide  wave-guide